OpenStudy (anonymous):

Find three real geometric means between 1/525 and 25/21

OpenStudy (amistre64):

if we write out the sequence like this g1 m1 m2 m3 g5

OpenStudy (amistre64):

then we can reconstruct it by knowing the formula for a geometric series

OpenStudy (amistre64):

gn = g1 * r^(n-1) gn/g1 = r^(n-1) (n-1)rt of (gn/g1) = r

OpenStudy (amistre64):

when the (n-1) part gets larger than 3 its prolly easier to use the exponential notation:\[\large r=\left(\frac{g_n}{g_1}\right)^{(\frac{1}{n-1})}\]

OpenStudy (anonymous):

ok how do plug these in

OpenStudy (amistre64):

define your gn and your g1 with the values given; since n=5 in this case, id use a 5 in place of n

OpenStudy (anonymous):

ok 5n

OpenStudy (anonymous):

whats g1

OpenStudy (amistre64):

g5 = 1/525 ; g1 = 25/21 or g1 = 1/525 ; g5 = 25/21 it doesnt matter which is which, as long as we are consistent thruout

OpenStudy (amistre64):

if we pick the first option: g5 = 1/525 ; g1 = 25/21 .... i get an r value of 5

OpenStudy (amistre64):

g2 = g1 * r^1 g3 = g1 * r^2 g4 = g1 * r^3 which then would give us g2 = 25/21 * 5^1 g3 = 25/21 * 5^2 g4 = 25/21 * 5^3

OpenStudy (anonymous):

so do we multiply

OpenStudy (amistre64):

make that r = 1/5 if we go that route ....

OpenStudy (amistre64):

yes, we multiply after we set it all up. im sure that there is a simpler way to go about it, but i never remember it ....

OpenStudy (anonymous):

92=5.95

OpenStudy (anonymous):

g2=29.76

OpenStudy (anonymous):

g4=148.8

OpenStudy (amistre64):

i made a mistake in the "r" with my setup; try using this instead g2 = 25/21 * 1/5^1 g3 = 25/21 * 1/5^2 g4 = 25/21 * 1/5^3

OpenStudy (anonymous):

isn't it the same

OpenStudy (anonymous):

oh no ok

OpenStudy (amistre64):

no, multipling stuff by "5" is not the same as multiplying stuff by "1/5"

OpenStudy (anonymous):

g2=.2380 g3=.0476 g4=.0095

OpenStudy (amistre64):

that looks better to me; just be sure they want it in decimal form before submitting; else keep it in fractions

OpenStudy (anonymous):

ok thanks